In a triangle ABC, angle B is 5 times angle A and angle C is 6(degree's) less than 4 times angle A?
Answer by algebrahouse.com(1659)

"In a triangle ABC, angle B is 5 times angle A and angle C is 6(degree's) less than 4 times angle A?"

x = angle A
5x = angle B {angle B is five times angle A}
4x - 6 = angle C {angle C is six less than four times A}

Angles of a triangle add up to 180°

x + 5x + 4x - 6 = 180 {angles add up to 180°}
10x - 6 = 180 {combined like terms}
10x = 186 {added 6 to both sides}
x = 18.6 {divided both sides by 10}
5x = 93 {substituted 18.6, in for x, into 5x}
4x - 6 = 68.4 {substituted 18.6, in for x, into 4x - 6}

the three angles are 18.6°, 93°, and 68.4°

For more help from me, visit: www.algebrahouse.com